Can someone help me with this solution for this.
One card is selected at random from a standard 52-card deck of playing cards. Find the probability that the card selected is a jack.

Answer by checkley77(12844)

p=success/total possibiliyies
p=4/52=1/13=.0769 or 7.69% chance of selecting a JACK.